OpenStudy (anonymous):

I am trying to simplify this : 2(x-1)(x+1)^2+2(x+1)(x-1)^2

OpenStudy (anonymous):

\[4(x-1)^{2}\]

myininaya (myininaya):

2(x-1)(x+1)[(x+1)+(x-1)] 2(x-1)(x+1)(2x)

OpenStudy (anonymous):

but where do the squares go

myininaya (myininaya):

mathstud did you ever look at my attachment about this problem i already simplied it for you

OpenStudy (anonymous):

i just dont understand you got rid of the squares

myininaya (myininaya):

i factored out (x+1) and (x-1)

myininaya (myininaya):

i also factored out a 2

OpenStudy (anonymous):

can you post that doc again? please

OpenStudy (anonymous):

4x(x-1)^2

OpenStudy (anonymous):

thanks

myininaya (myininaya):

2x^2y+2xy^2 we can factor this right? both terms have a 2, and a x, and a y in common so we have 2x^2y+2xy^2=2xy(x+y)

OpenStudy (anonymous):

(x-1)(x+1) = (x^2-1) can be factored.

myininaya (myininaya):

yes x^2-1 can be factorer its factors are x-1 and x+1 gj

OpenStudy (anonymous):

u get it, mathstud?

OpenStudy (anonymous):

I mean, do you see where the squares go?

OpenStudy (anonymous):

Gone away to study it...

OpenStudy (anonymous):

hey, why is there a max at 1 and a min 0

OpenStudy (anonymous):

I got x=1, x=-1, and x=0

myininaya (myininaya):

i got a max at x=0 and a min at x=1 and x=-1

OpenStudy (anonymous):

It's 4x(x-1)^2 so x=0 and 1(repeated).

myininaya (myininaya):

estudier its suppose to be 4x(x^2-1) or 4x(x-1)(x+1)

OpenStudy (anonymous):

Oops, sorry...(beg for forgiveness)

OpenStudy (anonymous):

0 and +-1 looks good...